what is the equation of a line that is parallel to the line y=-3x-4 and passes through the point (-4,6)
Answer by lwsshak3(11628)

what is the equation of a line that is parallel to the line y=-3x-4 and passes through the point (-4,6)
**
equation of a line: y=mx+bd, m=slope, b=y-intercept
y=-3x-4
m=-3
a parallel line will have the same slope
y=-3x+b
solve for b using coordinates of given point (-4,6)
6=-3*-4+b
b=-12+6=-6
Equation of parallel line: y=-3x-6